OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. 25x2 − 40x − 16 9a2 − 20a − 25 25b2 − 15b + 9 16x2 − 56x + 49

OpenStudy (triciaal):

a perfect square trinomial (a + b)^2 = a^2 + 2 a b + b^2 rephrasing the question ,which expression is in this format? hint: when you square an entity it is always positive

OpenStudy (anonymous):

16x2 − 56x + 49

OpenStudy (anonymous):

\[16 x^2-56 x+49=(4 x-7)^2 \]